# Feature Hedging: Correlated Features Break Narrow Sparse Autoencoders

### v1

@@ Abstract
It is assumed that sparse autoencoders (SAEs) decompose polysemantic activations into interpretable linear directions, as long as the activations are composed of sparse linear combinations of underlying features. However, we find that if an SAE is more narrow than the number of underlying “true features” on which it is trained, and there is correlation between features, the SAE will merge components of correlated features together, thus destroying monosemanticity. In LLM SAEs, these two conditions are almost certainly true. This phenomenon, which we call feature hedging, is caused by SAE reconstruction loss, and is more severe the narrower the SAE. In this work, we introduce the problem of feature hedging and study it both theoretically in toy models and empirically in SAEs trained on LLMs. We suspect that feature hedging may be one of the core reasons that SAEs consistently underperform supervised baselines. Finally, we use our understanding of feature hedging to propose an improved variant of matryoshka SAEs. Our work shows there remain fundamental issues with SAEs, but we are hopeful that that highlighting feature hedging will catalyze future advances that allow SAEs to achieve their full potential of interpreting LLMs at scale.
@@ 1 Introduction
As large language models (LLMs) are deployed in real-world applications, it is increasingly important to understand their internal workings. Sparse autoencoders (SAEs) decompose the dense, polysemantic activations of LLMs into interpretable latent features [6, 2] using sparse dictionary learning [19]. SAEs have the advantage of operating completely unsupervised, and can easily be scaled to millions of neurons in its hidden layer (hereafter called “latents”) [22, 11].
@@
While SAEs showed promising results, recent work has cast doubt on the performance of SAEs relative to baseline techniques. Wu et al. [24] show that SAEs underperform on both concept steering and detection relative to baselines, and Kantamneni et al. [13] show that SAEs underperform simple linear probes on both in-domain and out-of-domain detection, even when the probes have very few training samples. The question, then, is why do SAEs underperform relative to other techniques? And if we can identify the problems holding back SAEs, can we then fix those problems?
@@
One fundamental issue with SAEs is the problem of feature absorption [5], where a more specific latent suppresses the firing a more general latent. For instance, an SAE may have a latent that appears to track “Cities in USA” but that arbitrarily fails to fire on the specific cities “New York” and “Detroit”, where a city-specific latent fires instead. Feature absorption requires underlying features to exist in a hierarchy, with a parent feature $f_{p}$ and a child feature $f_{c}$, where $f_{c}$ can only fire if $f_{p}$ is firing ($f_{c}\implies f_{p}$). Feature absorption is caused by SAE sparsity penalty, and becomes more severe the wider the SAE. An SAE encoder/decoder under feature absorption is shown in Figure 1(b).
@@
In this paper, we identify another fundamental issue with SAEs which we call feature hedging. In hedging, an SAE is too narrow to represent both features $f_{a}$ and $f_{b}$ with their own latents $l_{a}$ and $l_{b}$. Ideally, an SAE should assign a latent $l$ to either $f_{a}$ or $f_{b}$, and ignore the feature not being tracked. However, if $f_{a}$ and $f_{b}$ are either hierarchical as in absorption, or (anti-)correlated, then the SAE latent $l$ can reduce reconstruction error by incorrectly mixing in components of both $f_{a}$ and $f_{b}$. A sample SAE encoder and decoder experiencing hedging is shown in Figure 1(a). In an LLM SAE, hedging will look like each SAE latent has noise mixed into it, likely reducing the performance of the latent for both detection and steering. Unlike with absorption, hedging becomes worse the narrower the SAE: thus trying to reduce absorption by making the SAE narrower will simply result in more hedging instead. The differences between feature hedging and feature absorption are shown in Table 1.
@@
In LLM SAEs, the SAE is almost certainly narrower than the number of underlying features, as even extremely wide LLM SAEs appear to still miss features [22]. Furthermore, we should expect that nearly every feature in an LLM has some correlation to other features. We thus expect that hedging is the norm in SAEs trained on LLM activations and will distort the performance of LLM SAEs.
@@
Learns gerrymandered latents Mixes correlated features into latents
@@
(b) If the SAE is wide enough to track both features, we see feature absorption. The decoder for $l_{1}$ perfectly tracks $f_{1}$, but its encoder turns off if $f_{2}$ is also active. $l_{2}$ tracks $f_{2}$, but its decoder merges together both $f_{1}$ and $f_{2}$. This asymmetry between encoder and decoder is characteristic of feature absorption.
@@
A solution to feature absorption has been proposed in the form of matryoshka SAEs [4]. Matryoshka SAEs use nested SAE loss terms to enforce a hierarchy on the SAE latents, solving absorption by forcing the narrow inner levels of the SAE to reconstruct inputs on their own. However, as we show in this paper, matryoshka SAEs suffer more from hedging due to the inner matryoshka levels essentially being very narrow SAEs. Matryoshka SAEs thus trade off absorption for hedging.
In this work, we define and study feature hedging both theoretically in toy models and empirically in LLM SAEs. We show that hedging is worse the more narrow the SAE, and introduce a technique to characterize the amount of hedging present in a given SAE. We also study hedging and absorption in matryoshka SAEs, and show that it is possible to improve the monosemanticity of matryoshka SAEs by tuning the relative loss coefficients in each level of the matryoshka SAE to better balance the competing forces of absorption and hedging—though both problems remain present.
@@ 2 Background
An SAE decomposes an input activation $a\in\mathbb{R}^{D}$ into a hidden state $f$ consisting of $L$ hidden neurons, called “latents”. An SAE is composed of an encoder $W_{\text{enc}}\in\mathbb{R}^{L\times D}$, a decoder $W_{\text{dec}}\in\mathbb{R}^{D\times L}$, a decoder bias $b_{\text{dec}}\in\mathbb{R}^{D}$, and encoder bias $b_{\text{enc}}\in\mathbb{R}^{L}$, and a nonlinearity $\sigma$, typically ReLU or a variant like JumpReLU [20], TopK [11] or BatchTopK [3].
$f=$ $\sigma(W_{\text{enc}}(a-b_{\text{dec}})+b_{\text{enc}})$ (1)
$\hat{a}=$ $W_{\text{dec}}f+b_{\text{dec}}$ (2)
The SAE is trained with a reconstruction loss, typically Mean Squared Error (MSE), and a sparsity-inducing loss consisting of a function $\mathcal{S}$ that penalizes non-sparse representation with corresponding sparsity coefficient $\lambda$. For standard L1 SAEs, $\mathcal{S}$ is the L1 norm of $f$. For TopK and BatchTopK SAEs, there is no sparsity-inducing loss ($\mathcal{S}=0$) as the TopK function directly induces sparsity. There is sometimes also an additional auxiliary loss $\mathcal{L}_{aux}$ with coefficient $\alpha$ to ensure all latents fire. Standard L1 SAEs typically do not have an auxiliary loss [18]. The general SAE loss is
$\mathcal{L}=\|a-\hat{a}\|_{2}^{2}+\lambda\mathcal{S}+\alpha\mathcal{L}_{\text{aux}}.$ (3)
@@
A matryoshka SAE [4] extends the SAE definition by summing losses created by prefixes of SAE latents. This forces each sub-SAE to reconstruct input activations on its own, and incentivizes the SAE to place more common, general concepts into latents with smaller index number. A matryoshka SAE uses nested prefixes with sizes $\mathcal{M}=m_{1},m_{2},...m_{n}$ where $m_{1}<m_{2}<\ldots<m_{n}=L$, where $L$ is the number of latents in the full dictionary. Matryoshka SAE loss is:
$\mathcal{L}=\sum_{m\in\mathcal{M}}\left(\|a-\hat{a}_{m}\|_{2}^{2}+\lambda\mathcal{S}_{m}\right)+\alpha\mathcal{L}_{\text{aux}}$ (4)
Where $\hat{a}_{m}$ is the reconstruction for the SAE using the first $m$ latents, and $\mathcal{S}_{m}$ is the sparsity penalty applied to the first $m$ latents. For TopK and BatchTopK Matryoshka SAEs, there is no sparsity penalty ($\mathcal{S}_{m}=0$) as the TopK function directly imposes sparsity.
3 Studying hedging in single-latent SAEs
We begin by investigating hedging in the simplest possible toy SAE setting: an SAE with a single latent. We use a model with two true features $f_{1}$ and $f_{2}$. Each true feature $f$ is a random direction with unit-norm in $\mathbb{R}^{50}$, and $f_{1}\perp f_{2}$. Each feature fires with magnitude 1.0. Since we only have two features, an activation $a$ can consist of $a\in\{0,f_{1},f_{2},f_{1}+f_{2}\}$. There is no bias term added to the activations. Unless otherwise specified, $f_{1}$ fires with probability 0.25, and $f_{2}$ fires with probability 0.2. We use SAELens [1] to train a single-latent SAE on these activations.
@@ 3.1 Fully independent features
We first study the case when $f_{1}$ and $f_{2}$ fire independently. We find that the SAE correctly represents $f_{1}$ without any interference from $f_{2}$. However, the decoder bias has incorrectly learned to represent the direction of $f_{2}$, but with magnitude 0.2, equal to the probability of $f_{2}$ firing. The cosine similarities of the single SAE latent and SAE bias term with the true features is shown in Figure 2.
(a) The SAE encoder and decoder both correctly learn just $f_{1}$.
(b) The decoder bias incorrectly learns $f_{2}$.
Figure 2: Encoder, decoder, and decoder bias patterns for a toy model with 2 independent features.
We consistently find this pattern of the decoder bias merging in positive components of features not tracked by their own latent. In this sense, the decoder bias can be thought of as tracking an always-on feature, and thus is in a hierarchical relationship with every other feature of the model.
@@ 3.2 Hierarchical features
Next, we investigate what happens if $f_{1}$ and $f_{2}$ are in a hierarchy, so $f_{2}$ can only fire if $f_{1}$ fires, but $f_{1}$ can still fire on its own ($f_{2}\implies f_{1}$). We adjust the firing probability of $f_{2}$ so that $P(f_{2}|f_{1})=0.2$, and $P(f_{2}|\neg f_{1})=0$ (thus, $P(f_{2})=0.05$). In a two-latent SAE this setup would cause feature absorption. We plot the cosine similarities of our single latent with $f_{1}$ and $f_{2}$ in Figure 3(a).
(a) When features $f_{1}$ and $f_{2}$ form a hierarchy ($f_{2}\implies f_{1}$), the SAE incorrectly merges a component of $f_{2}$ into its single latent $l_{1}$.
(b) When features $f_{1}$ and $f_{2}$ are anti-correlated, the SAE incorrectly merges a negative component of $f_{2}$ into its single latent $l_{1}$.
Figure 3: Hedging occurs with hierarchical features or anti-correlated features.
Here we clearly see feature hedging. The single SAE latent has now merged in a component of $f_{2}$ into its single latent, so it’s now a mixture of $f_{1}$ and $f_{2}$. $f_{2}$ is merged roughly symmetrically into both the encoder and decoder of the SAE latent ($\cos(f_{2},l_{1})$ is about $\nicefrac{{1}}{{4}}$ of $\cos(f_{1},l_{1})$ in both encoder and decoder). This is unlike in feature absorption where there is an asymmetry in the encoder and decoder. This merging of features reduces the MSE loss of the SAE despite being a degenerate solution.
Increasing the L1 penalty of the SAE does not solve this problem. $f_{2}$ only fires if $f_{1}$ fires, so adding a positive component of $f_{2}$ into the encoder does not cause the latent to fire any more often.
@@ 3.3 Positively correlated features
Next, we change our setup so that $P(f_{2}|\neg f_{1})=0.1$ instead of 0. We still keep $P(f_{2}|f_{1})=0.2$, so that $f_{2}$ is more likely to fire if $f_{1}$ fires, but it can still fire on its own as well. The features are now merely correlated rather than following a strict hierarchy.
We now see hedging depending on the strength of the L1 penalty. When the L1 penalty is low, hedging is apparent. However, if the L1 penalty is high enough and the level of correlation is low enough, then the SAE will learn the correct features, as positive hedging increases the L0 of the SAE slightly relative to learning just $f_{1}$. Plots of the cosine similarity of the SAE encoder and decoder compared to true features are shown in Figure 4 with high and low sparsity penalties. If we use a full-width SAE, the SAE learns the true features despite the correlation (see Appendix A.1).
(a) Hedging still occurs with L1 coefficient of 0.001.
(b) No hedging occurs with L1 coefficient of 0.1.
Figure 4: Hedging occurs with positively-correlated features depending on the sparsity penalty.

### v2

@@ Abstract
It is assumed that sparse autoencoders (SAEs) decompose polysemantic activations into interpretable linear directions, as long as the activations are composed of sparse linear combinations of underlying features. However, we find that if an SAE is more narrow than the number of underlying “true features” on which it is trained, and there is correlation between features, the SAE will merge components of correlated features together, thus destroying monosemanticity. In LLM SAEs, these two conditions are almost certainly true. This phenomenon, which we call feature hedging, is caused by SAE reconstruction loss, and is more severe the narrower the SAE. In this work, we introduce the problem of feature hedging and study it both theoretically in toy models and empirically in SAEs trained on LLMs. We suspect that feature hedging may be one of the core reasons that SAEs consistently underperform supervised baselines. Finally, we use our understanding of feature hedging to propose an improved variant of matryoshka SAEs. Importantly, our work shows that SAE width is not a neutral hyperparameter: narrower SAEs suffer more from hedging than wider SAEs.
@@ 1 Introduction
As large language models (LLMs) are deployed in real-world applications, it is increasingly important to understand their internal workings. Sparse autoencoders (SAEs) decompose the dense, polysemantic activations of LLMs into interpretable latent features (Cunningham et al., 2024; Bricken et al., 2023) using sparse dictionary learning (Olshausen & Field, 1997). SAEs have the advantage of operating completely unsupervised, and can easily be scaled to millions of neurons in its hidden layer (hereafter called “latents”  [^1] We use the term “latents” for the hidden neurons of the SAE to avoid overloading the term “feature”. We use “feature” only to describe interpretable concepts represented by the model.)(Templeton et al., 2024; Gao et al., 2024).
@@
While SAEs showed promising results, recent work has cast doubt on the performance of SAEs relative to baseline techniques. Wu et al. (2025) show that SAEs underperform on both concept steering and detection relative to baselines, and Kantamneni et al. (2025) show that SAEs underperform simple linear probes on both in-domain and out-of-domain detection, even when the probes have very few training samples. The question, then, is why do SAEs underperform relative to other techniques? And if we can identify the problems holding back SAEs, can we then fix those problems?
@@
One fundamental issue with SAEs is the problem of feature absorption (Chanin et al., 2024), where a more specific latent suppresses the firing a more general latent. For instance, an SAE may have a latent that appears to track “Cities in USA” but that arbitrarily fails to fire on the specific cities “New York” and “Detroit”, where a city-specific latent fires instead. Feature absorption requires underlying features to exist in a hierarchy, with a parent feature $f_{p}$ and a child feature $f_{c}$, where $f_{c}$ can only fire if $f_{p}$ is firing ($f_{c}\implies f_{p}$). Feature absorption is caused by SAE sparsity penalty, and becomes more severe the wider the SAE. An SAE encoder/decoder under feature absorption is shown in Figure 1(b).
@@
In this paper, we identify another fundamental issue with SAEs which we call feature hedging. In hedging, an SAE is too narrow to represent both features $f_{a}$ and $f_{b}$ with their own latents $l_{a}$ and $l_{b}$. Ideally, an SAE should assign a latent $l$ to either $f_{a}$ or $f_{b}$, and ignore the feature not being tracked. However, if $f_{a}$ and $f_{b}$ are either hierarchical as in absorption, or (anti-)correlated, then the SAE latent $l$ can reduce reconstruction error by incorrectly mixing in components of both $f_{a}$ and $f_{b}$. A sample SAE encoder and decoder experiencing hedging is shown in Figure 1(a). In an LLM SAE, hedging will look like each SAE latent has noise mixed into it, reducing the performance of the latent for both detection and steering. Unlike with absorption, hedging becomes worse the narrower the SAE: thus trying to reduce absorption by making the SAE narrower will simply result in more hedging instead. The differences between hedging and absorption are shown in Table 1.
@@
In LLM SAEs, the SAE is almost certainly narrower than the number of underlying features, as even extremely wide LLM SAEs appear to miss features (Templeton et al., 2024). Furthermore, we expect that nearly every feature in an LLM has positive and negative correlations to many features. We thus expect that hedging is the norm in LLM SAEs and will significantly distort their performance.
@@
Learns gerrymandered latents Learns polysemantic mixtures of features
@@
(b) Adding a new latent to the SAE so it is wide enough to track both features, we see feature absorption. The decoder for $l_{1}$ perfectly tracks $f_{1}$, but its encoder turns off if $f_{2}$ is also active. $l_{2}$ tracks $f_{2}$, but its decoder mixes $f_{1}$ and $f_{2}$. Asymmetry between encoder and decoder is characteristic of absorption.
@@
A solution to feature absorption has been proposed in the form of matryoshka SAEs (Bussmann et al., 2025). Matryoshka SAEs use nested SAE loss terms to enforce a hierarchy on the SAE latents, solving absorption by forcing the narrow inner levels of the SAE to reconstruct inputs on their own. However, as we show in this paper, matryoshka SAEs suffer more from hedging due to the inner matryoshka levels essentially being very narrow SAEs. Matryoshka SAEs thus trade off absorption for hedging.
In this work, we define and study feature hedging both theoretically in toy models and empirically in LLM SAEs. We show that hedging is worse the more narrow the SAE, and introduce a technique to characterize the amount of hedging present in a given SAE. We also study hedging and absorption in matryoshka SAEs, and show that it is possible to improve the monosemanticity of matryoshka SAEs by tuning the relative loss coefficients in each level of the matryoshka SAE to better balance the competing forces of absorption and hedging—though both problems remain present. We show as well that SAE width is not a neutral hyperparameter: narrow SAEs suffer more from hedging than wider SAEs.
@@ 2 Background
An SAE decomposes an input activation $x\in\mathbb{R}^{D}$ into a hidden state $f$ consisting of $L$ hidden neurons, called “latents”. An SAE is composed of an encoder $W_{\text{enc}}\in\mathbb{R}^{L\times D}$, a decoder $W_{\text{dec}}\in\mathbb{R}^{D\times L}$, a decoder bias $b_{\text{dec}}\in\mathbb{R}^{D}$, and encoder bias $b_{\text{enc}}\in\mathbb{R}^{L}$, and a nonlinearity $\sigma$, typically ReLU or a variant like JumpReLU (Rajamanoharan et al., 2024), TopK (Gao et al., 2024) or BatchTopK (Bussmann et al., 2024).
$z=$ $\sigma(W_{\text{enc}}(x-b_{\text{dec}})+b_{\text{enc}})$ (1)
$\hat{x}$= $W_{\text{dec}}z+b_{\text{dec}}$ (2)
The SAE is trained with a reconstruction loss, typically Mean Squared Error (MSE), and a sparsity-inducing loss consisting of a function $\mathcal{S}$ that penalizes non-sparse representation with corresponding sparsity coefficient $\lambda$. For standard L1 SAEs, $\mathcal{S}$ is the L1 norm of $f$. For TopK and BatchTopK SAEs, there is no sparsity-inducing loss ($\mathcal{S}=0$) as the TopK function directly induces sparsity. There is sometimes also an additional auxiliary loss $\mathcal{L}_{aux}$ with coefficient $\alpha$ to ensure all latents fire. Standard L1 SAEs typically do not have an auxiliary loss (Olah et al., 2024). The general SAE loss is
$\mathcal{L}=\|x-\hat{x}\|_{2}^{2}+\lambda\mathcal{S}+\alpha\mathcal{L}_{\text{aux}}.$ (3)
@@
A matryoshka SAE (Bussmann et al., 2025) extends the SAE definition by summing losses created by prefixes of SAE latents. This forces each sub-SAE to reconstruct input activations on its own, and incentivizes the SAE to place more common, general concepts into latents with smaller index number. A matryoshka SAE uses nested prefixes with sizes $\mathcal{M}=m_{1},m_{2},...m_{n}$ where $m_{1}<m_{2}<\ldots<m_{n}=L$, where $L$ is the number of latents in the full dictionary. Matryoshka SAE loss is:
$\mathcal{L}=\sum_{m\in\mathcal{M}}\left(\|x-\hat{x}_{m}\|_{2}^{2}+\lambda\mathcal{S}_{m}\right)+\alpha\mathcal{L}_{\text{aux}}$ (4)
Where $\hat{x}_{m}$ is the reconstruction for the SAE using the first $m$ latents, and $\mathcal{S}_{m}$ is the sparsity penalty applied to the first $m$ latents. For TopK and BatchTopK Matryoshka SAEs, there is no sparsity penalty ($\mathcal{S}_{m}=0$) as the TopK function directly imposes sparsity.
3 Toy models of feature hedging
The linear representation hypothesis (LRH) states that features in LLMs are represented as nearly-orthogonal linear directions in representation space (Bricken et al., 2023). The goal of SAEs, then, is to recover these underlying “true features” of the model, where each latent of the SAE decoder perfectly matches an underlying feature of the model. In real LLMs we do not have ground-truth knowledge of these underlying features, making it difficult to know if SAEs are succeeding at recovering model features. Fortunately, it is easy to create synthetic training data for SAEs that follows the LRH and gives us ground-truth knowledge of the underlying features. This allows us to understand when SAEs will learn the underlying features of the model, and when SAEs fail.
We define a toy model consisting of $N$ true features $F\in\mathbb{R}^{N\times D}$, where each $\|f_{i}\rVert_{2}=1$. These features are mutually orthogonal, so $\forall i\neq j,f_{i}\cdot f_{j}=0$. Each feature $f_{i}$ has a corresponding firing probability $p_{i}\in[0,1]$. For each sample, we generate a binary activation vector $\mathbf{a}\in\{0,1\}^{N}$ where $a_{i}\sim\text{Bernoulli}(p_{i})$ indicates whether feature $f_{i}$ is active (fires). The model can incorporate feature dependencies by conditioning firing probabilities: $a_{i}|\mathbf{a}_{-i}\sim\text{Bernoulli}(p_{i}(\mathbf{a}_{-i}))$, where $\mathbf{a}_{-i}$ denotes the activation states of all other features. We then generate SAE training samples $x$ from this model as $x=\sum_{i=1}^{N}a_{i}f_{i}$.
We say that an SAE is correct or monosemantic for this toy model if every latent in the SAE dictionary matches a true feature direction, and each SAE latent corresponds to a different true feature. Formally, there exists a bijection $\pi:\{1,\ldots,L\}\to\{1,\ldots,N\}$ such that $\cos(W_{\text{dec},i},f_{\pi(i)})=1$ for all $i\in\{1,\ldots,L\}$. We only investigate SAEs where $L\leq N$ in our toy experiments. We say an SAE is polysemantic if some SAE latents contain positive or negative components of multiple true features, so there exists at least one latent $i\in\{1,\ldots,L\}$ such that $|\{j\in\{1,\ldots,N\}:|W_{\text{dec},i}\cdot f_{j}|>\epsilon\}|>1$ for some threshold $\epsilon>0$.
For all SAEs in this section, we train on 15M synthetic activations using SAELens (Bloom et al., 2024). In this section we show plots of the cosine similarity between the SAE encoder / decoder and the true features. Each cell $i,j$ in these plots is simply $\cos(W_{\text{enc},i}^{T},f_{j})$ and $\cos(W_{\text{dec},i},f_{j})$, respectively. We re-arrange the indices of the SAE latents to best align visually with true features.
@@ 3.1 Fully independent features
We first study the case of a toy model with $N=4$ independent features. Features 1-3 fire with probability $0.25$, and feature 4 fires with probability $0.2$. We plot the encoder / decoder cosine similarity with true features in Figure 2. When features fire independently, the SAE learns correct features regardless of the width of the SAE.
Figure 2: SAE with three latents (left) and four latents (right) trained on a toy model with independent features. Both SAEs learn correct features.
Unfortunately, real LLMs do not have fully independent features. SAEs were first studied under toy models with independent features (Elhage et al., 2022), and this is likely why the field was not aware of feature hedging earlier.
@@ 3.2 Hierarchical features
Next, we explore true features that fire hierarchically. We modify the toy model from Section 3.1 above, and set $f_{3}$ as the parent feature and $f_{4}$ as the child feature, so $f_{4}\implies f_{3}$. That is, $f_{4}$ cannot fire unless $f_{3}$ is also firing. Hierarchical features cause feature absorption in SAEs that are wide enough to represent both the parent and child feature, but what happens if the SAE is not wide enough to represent the child latent? This is important as this is the intuition behind why Matryoshka SAEs work to combat absorption: if inner SAE levels are too narrow to represent both parent and child features, we hope that only the parent will be represented. We show results in Figure 3.
Figure 3: SAEs trained on a toy model with hierarchical features ($f_{3}\implies f_{4}$). When the SAE is too narrow to represent $f_{4}$ (left), we see hedging where latent 3 mixes . When the SAE is wide enough to contain both $f_{3}$ and $f_{4}$ (right), we see feature absorption.
As expected, in the full-width SAE we see a classic feature absorption pattern. The parent latent encoder, $l_{3}$, learns $\lnot f_{4}\land f_{3}$, disabling the latent from firing if $f_{4}$ is present. The child latent, $l_{4}$, mixes both $f_{3}$ and $f_{4}$ together in the decoder.
However, when the SAE is too narrow to represent $f_{4}$, we now see that latent $l_{3}$ mixes a component of $f_{4}$ into both its encoder and decoder! We refer to this as feature hedging. The SAE is learning an incorrect, polysemantic latent that mixes correlated features.
While we expect this will be a problem for all SAEs, it is particularly problematic for Matryoshka SAEs, as Matryoshka SAEs combat absorption by using inner SAE levels that are too narrow to contain both parent and child features. However, as we see here, this causes hedging.
We show that MSE loss directly causes hedging with hierarchical features in Appendix A.1.
@@ 3.3 Positively correlated features
Hierarchy is a particularly extreme form of positive correlation, where a feature can only fire if another feature also fires. Next, we relax that restriction, and investigate what happens if a feature is merely more likely to fire along with another feature, but can still fire on its own as well. We modify the toy model so that $p_{4}=0.2$ if $f_{3}$ fires, but $p_{4}=0.1$ if $f_{3}$ does not fire, adding a small positive correlation between $f_{3}$ and $f_{4}$. We show results Figure 4
Figure 4: SAEs trained on a toy model with positive correlation between features $f_{3}$ and $f_{4}$. When the SAE is too narrow to represent $f_{4}$ (left), we still see hedging in latent 3. When the SAE is wide enough to contain both $f_{3}$ and $f_{4}$ (right), the SAE learns correct features.
We still see that the SAE is mixing a positive component of $f_{4}$ into $l_{3}$, despite there no longer being perfect hierarchy! We also see that if we extend the SAE width so that $f_{4}$ is tracked by its own latent $l_{4}$, there is now no absorption at all, as absorption requires (nearly) hierarchical features to arise.